If $1,000 is deposited for 6 years in a savings account earning 3.6% interest compounded monthly, how much interest is earned during the sixth year?
Answer by Theo(13342)

the formula to use is f = p * (1 + r) ^ n.

f is the future value
p is the present value
r is the interest rate per time period
n is the number of time periods

the investment period is 6 years * 12 = 72 months.
the interest rate per month is 3.6% / 100 = .036 per year / 12 = .003 per month.

the formula becomes f = 1000 * (1 + .003) ^ 72 = 1240.701129.

that would be at the end of the sixth year.

at the end of the fifth year, the fomula becomes f = 1000 * (1 + .003) ^ 60.

solve for f to get f = 1196.894803.

subtract 1196.894803 from 1240.701129 and you get 43.80632619.

that's the interest earned during the sixth year.

this can be seen more clearly in the following spreadsheet printout.

$$$

$$$

the column headings are:

eom - end of month
rembal - remaining balance
int - interest earned from remaining balance in the previous month.
cumint - cumulative sum of interest earned up to the end of month indicated.